Solve the following equation for A : 2A/3 = 8 + 4A
Answer by checkley75(3666)

2A/3=8+4A NOW CROSS MULTIPLY
2A=3(8+4A)
2A=24+12A
2A-12A=24
-10A=24
A=24/-10
A=-2.4
PROOF
2*-2.4/3=8+4*-2.4
-4.8/3=8-9.6
-1.6=-1.6